A straight line makes an intercept on the y-axis twice as long as that on the x-axis and is at a unit distance from the origin. Determine its equation.

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\y=mx+c \\=>y_{interc}=c \\=>x_{interc}=-c/m \\=>m=-1 \\distance={c \over \sqrt{1+m^2}} \\={c \over \sqrt{2}}=\pm 1 \\=>c=\pm \sqrt{2} \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty